# Expansion Span: Combining Fading Memory and Retrieval in Hybrid State Space Models

### v1

@@ 3.1 Attention
At the heart of modern LLMs is the Attention mechanism, whose role is to construct a contextualized representation of their input. The input to Attention, $x\in\mathbb{R}^{L\times d}$, is a tensor of $L$ tokens, each with dimension $d$. Attention is parameterized by $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d_{\textrm{model}}}$ and $W_{o}\in\mathbb{R}^{d_{\textrm{model}}\times d}$, which are used to construct linear projections of the input: $Q=xW_{Q}$, $K=xW_{K}$, $V=xW_{V}\in\mathbb{R}^{L\times d_{\textrm{model}}}$. After adding positional information to the keys and queries, the output of the Attention layer is [^1] For simplicity, we consider Single-Head Attention in this exposition. $Attention(Q,K,V)=\left[\mathrm{softmax}\left(\frac{QK^{T}}{\sqrt{d_{\textrm{model}}}}\right)V\right]W_{o}\in\mathbb{R}^{L\times d}$. The realization of the attention score matrix, $QK^{T}\in\mathbb{R}^{L\times L}$, grows quadratically in the sequence length, $L$. Moreover, the output of Attention is typically followed by a feed-forward layer (FFN), which expands and contracts the dimension of the input. While these FFNs are generally regarded as being the primary computational bottlenecks of Attention-based models, when the sequence length exceeds the expanded dimension, Attention becomes the primary bottleneck. Since LLMs consist of many layers of Attention, this computation is particularly expensive during training, where activations are cached. To make the Attention mechanism computationally amenable to longer contexts during training, we draw inspiration from RAG and Sparse Attention methods.
@@ 3.2 Amnesic Attention
The crux of our method is in the chunking of the $L$ tokens in the input sequence $x$, into chunks of size $M$. Namely, we begin by computing projections $Q=xW_{Q}$, $K=xW_{K}$, $V=xW_{V}$ and adding positional information as in standard Attention. Next, each of the $Q,K,V\in\mathbb{R}^{L\times d_{\textrm{model}}}$ projections are split into $T=\frac{L}{M}$ chunks of $M$ tokens in each along the sequence dimension, yielding $Q_{i},K_{i},V_{i}\in\mathbb{R}^{M\times d_{\textrm{model}}}$ for $i=1,\ldots,T$. A naive way to reduce the quadratic cost of Attention is by applying it independently on each of these chunks, $A_{i}=Attention(Q_{i},K_{i},V_{i})$ and then concatenate and project them to get the final output, $\textrm{Concatenate}(A_{1},A_{2},\ldots,A_{T})$. However, since the chunks are processed independently, there is no information exchanged between them. Hence, this naive—though efficient—Attention mechanism, which we refer to as “SE-Attn-NoMem”, cannot model time dependencies on contexts larger than the chunk size.
@@ 3.3 Eidetic Retrieval Attention
In this section, we improve upon SE-Attn-NoMem by allowing different chunks to exchange information while minimizing compute; we call this Attention mechanism SE-Attn. To this end, we augment the processing of each chunk with a mechanism to retrieve tokens from previous chunks. In particular, we allow chunk $i$ to retrieve tokens from chunks $11,2,\ldots,i-$ 1 and, when the most relevant chunks are selected, append their tokens to its context (its “expansion span”). SE-Attn takes as input a sequence, $x\in\mathbb{R}^{L\times d}$ and computes projections $Q=xW_{Q}$, $K=xW_{K}$, $V=xW_{V}$ followed by the addition of RoPE [29] embeddings as in standard Attention. As described in Sec. 3.2, $Q,K,V$ are split into $T$ chunks with $M$ tokens in each, yielding tuples $(Q_{i},K_{i},V_{i})$, $i=1,\ldots,T$. Additionally, $Q,K,V$ are split into a second set of $U$ chunks with $S$ tokens in each, yielding tuples $(Q^{\textrm{Mem}}_{j},K^{\textrm{Mem}}_{j},V^{\textrm{Mem}}_{j})$ where $Q^{\textrm{Mem}}_{j},K^{\textrm{Mem}}_{j},V^{\textrm{Mem}}_{j}\in\mathbb{R}^{S\times d_{\textrm{model}}}$ for $j=1,\ldots,U$. We refer to these tuples as “memory blocks”; in SE-Attn, the query from each chunk, $Q_{i}$, attends not only to $K_{i}$, but also to a set of retrieved memory blocks which populate SE-Attn’s expansion span. In particular, each $Q_{i}$ retrieves $k$ (top-$k$) key/value memory blocks from the past [^2] By “the past,” we mean tokens in the sequence that came before tokens in the chunk.: $\left(K^{\textrm{Mem}}_{\phi_{i}(U)_{1}},V^{\textrm{Mem}}_{\phi_{i}(U)_{1}}\right),\ldots,\left(K^{\textrm{Mem}}_{\phi_{i}(U)_{k}},V^{\textrm{Mem}}_{\phi_{i}(U)_{k}}\right)$ where $\phi_{i}(U)_{j}$ denotes the index of the $j$-th memory block selected by the $i$-th chunk. These retrieved blocks are then appended to the chunks’ keys and values, and SE-Attn computes the Attention output for the $i$th chunk as in Equation 3.
@@
$A_{i}^{\textrm{SE-Attn}}$ $=Attention(Q_{i},\tilde{K}_{i},\tilde{V}_{i})$ (3)
Afterwards, each chunk’s Attention outputs are concatenated and projected to obtain the SE-Attn layer output:
$o^{\textrm{SE-Attn}}$ $=\textrm{Concatenate}(A_{1}^{\textrm{SE-Attn}},A_{2}^{\textrm{SE-Attn}},\dots,A_{T}^{\textrm{SE-Attn}}).$ (4)
Note that Equation 3 is a form of Cross-Attention since we are not concatenating memory query tokens to the chunk’s query. This is to preserve causality, as the retrieved tokens cannot attend to the chunk’s tokens. Next, we discuss how the memory tokens are retrieved.
Memory Retrieval. Each chunk $x_{i}$ must judiciously select which memory blocks to retrieve from the past. To do this efficiently, we associate a 1-dimensional tensor, $c_{j}\in\mathbb{R}^{d_{\textrm{model}}}$, to each of the $j=1,\ldots,U$ memory blocks which act as a compressed representation of each memory block. To determine which memory blocks $Q_{i}$ should attend to, we compute a “relevancy score” between $Q_{i}$ and each $c_{j}$, which measures how relevant memory block $j$ is to chunk $i$. This relevancy score is implemented with a Cross-Attention score between chunks and compressed memory block representations. More specifically, recalling that $Q_{i}\in\mathbb{R}^{M\times d_{\textrm{model}}}$, we compute relevancy score $R_{ij}\in\mathbb{R}$ between chunk $i\in\{1,2,\ldots,T\}$ and memory block $j\in\{1,2,\ldots,U\}$ as follows:
$R_{ij}$ $=\sum_{t=1}^{M}(Q_{i}c_{j})_{t}.$ (5)
$R_{i}\in\mathbb{R}^{U}$ represents the relevancy score between chunk $i$ and all memory blocks. However, since chunk $i$ should only retrieve memory blocks that came before it temporally, we add a mask to $R_{i}$ and then apply Softmax to obtain the final scores, $\tilde{R}_{i}$, between chunk $i$ and all memory blocks as follows:
@@
Compressed Memory Blocks. Next, we discuss how we construct the compressed memory block representations, $c_{j}$. Recently, Landmark Attention [23] considered using “landmark” tokens to obtain compressed representations of memory blocks. We consider using landmark tokens to construct $c_{j}$ in Sec. 8. In SE-Attn, we consider a simpler approach which we found to work well. Namely, for each memory block, $(Q^{\textrm{Mem}}_{j},K^{\textrm{Mem}}_{j},V^{\textrm{Mem}}_{j})$, we perform standard non-causal Attention, $A^{\textrm{Mem}}_{j}=\mathrm{softmax}\left(\frac{Q^{\textrm{Mem}}_{j}(K^{\textrm{Mem}}_{j})^{T}}{\sqrt{d_{\textrm{model}}}}\right)V^{\textrm{Mem}}_{j}\in\mathbb{R}^{S\times d_{\textrm{model}}}$; we consider non-causal Attention since we are interested in a global representation where all tokens within the memory block can attend to each other. Next, we simply compute the mean of these weighted memory tokens as the compressed representation as:
$c_{j}$ $=\frac{1}{S}\sum_{t=1}^{S}(A^{\textrm{Mem}}_{j})_{t}\in\mathbb{R}^{d_{\textrm{model}}}$ (7)
where $(A^{\textrm{Mem}}_{j})_{t}$ denotes the $t$th row of $A^{\textrm{Mem}}_{j}$.
@@ 3.4 Training SE-Attn with LoRA
In this paper, we fine-tune models pre-trained with standard Attention; however, we fine-tune them with SE-Attn, which modifies the pre-trained Attention mechanism by introducing a retrieval mechanism. SE-Attn repurposes the model’s Attention parameters $(W_{Q},W_{K},W_{V},W_{o})$ to perform retrieval. In order to efficiently train the model to learn to use SE-Attn, we use a variant of LoRA. Recently, [3] introduced LoRA+, a variant of LoRA [17] designed to fine-tune Transformer models on long contexts. LoRA fine-tunes Attention parameters with low rank adapter matrices. LoRA+ differs from LoRA by also training embedding and normalization layers.
Recently, [10] found that pure SSMs can be fine-tuned by training the SSM projection layers with LoRA. Since we fine-tune on long contexts, we prioritize efficient training, and consequently apply LoRA only to the Attention layers of our hybrid models. However, it is common for SSM layers to include a 1D convolution layer after their initial projection in order to perform sequence mixing [38, 12, 22, 7, 8]. The 1D convolution parameters constitute a very small portion of the model parameters ($\sim$0.7for Mamba-2-Hybrid 2.7B), but as discussed further in Sec. 4.6, we found that training these 1D convolution layers in conjunction with LoRA+ improved our models’ performance on long-context tasks. We refer to this augmented LoRA+ variation as HyLoRA and use it to fine-tune all of our hybrid models.
@@ 4.1 Experimental Setup
Models. We enhance the recall capabilities of pre-trained hybrid SSM models by fine-tuning them with different Attention layers on spans of tokens longer than the ones used for pre-training. We consider Mamba-2-Hybrid 2.7B [^3] We use the Mamba-2-Hybrid weights available on HuggingFace: https://huggingface.co/state-spaces/mamba2attn-2.7b. [34] as our representative SSM hybrid model. We also explore expanding the span of Transformer models, and consider Llama1 7B [31] as our representative model. Both pre-trained model checkpoints were obtained from HuggingFace and were pre-trained with a context size of 2048. Throughout our experiments, “Non-fine-tuned” refers to the pre-trained model with no fine-tuning.
@@
Datasets. We fine-tune models using a common language dataset and provide additional results when fine-tuning on PG-19 [26] and a mixture of language and code dataset in Sec. 7.
@@
Baselines. For Mamba-2-Hybrid, we compare SE-Attn to three Attention variants: standard full Attention (“Full-Attn”) [33], Sliding Window Attention (“SW-Attn”) [1], and Shifted Sparse Attention ($S^{2}$-Attn) [3]. The Full-Attn baseline serves as the paragon, as the other methods aim to approximate it. For Llama1, due to computational constraints, we only compare against SW-Attn and $S^{2}$-Attn. All Attention implementations are based on FlashAttention-2 [5].
@@
Training Procedure. We adopt the training recipe used in [3] to fine-tune our models, with the exception of using HyLoRA for Mamba-2-Hybrid. Additional training details are provided in Sec. 9.
@@
Evaluation Metrics. We assess the performance of our models across a range of common benchmarks. To assess the predictive capabilities of our fine-tuned models on long sequence lengths, we measure perplexity (PPL) on the PG-19 validation set. To assess performance on more real-world language tasks in the short and long-context settings, we evaluate our models on various LM Evaluation Harness tasks [11]. To measure our models’ performance on long-context tasks, we evaluate on the challenging RULER [16] benchmark. We consider eleven RULER tasks, encompassing needle-in-a-haystack, variable-hopping, and aggregation tasks; we aggregate these metrics into five groups as explained in Sec. 11.4.
4.2 Evaluating with Full Attention

### v2

@@ 3.1 Attention
At the heart of modern LLMs is the Attention mechanism, whose role is to construct a contextualized representation of its input. The input to Attention, $x\in\mathbb{R}^{L\times d}$, is a tensor of $L$ tokens, each with dimension $d$. Attention is parameterized by $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d_{\textrm{model}}}$ and $W_{o}\in\mathbb{R}^{d_{\textrm{model}}\times d}$, which are used to construct linear projections of the input: $Q=xW_{Q}$, $K=xW_{K}$, $V=xW_{V}\in\mathbb{R}^{L\times d_{\textrm{model}}}$. After adding positional information to the keys and queries, the output of the Attention layer is [^1] For simplicity, we consider Single-Head Attention in this exposition. $\textrm{Attention}(Q,K,V)=\left[\mathrm{softmax}\left(\frac{QK^{T}}{\sqrt{d_{\textrm{model}}}}\right)V\right]W_{o}\in\mathbb{R}^{L\times d}$. The realization of the attention score matrix, $QK^{T}\in\mathbb{R}^{L\times L}$, grows quadratically in the sequence length, $L$. Moreover, the output of Attention is typically followed by a feed-forward layer (FFN), which expands and contracts the dimension of the input. While FFNs are generally regarded as being the primary computational bottlenecks of Attention-based models for short contexts, when the sequence length exceeds the expanded dimension, Attention becomes the primary bottleneck. Since LLMs consist of many layers of Attention, this computation is particularly expensive during training, where activations are cached. To make the Attention mechanism computationally amenable to longer contexts during training, we draw inspiration from RAG and Sparse Attention methods.
@@ 3.2 Amnesic Attention
The crux of our method is in the chunking of the $L$ tokens in the input sequence $x$, into chunks of size $M$. Namely, we begin by computing projections $Q=xW_{Q}$, $K=xW_{K}$, $V=xW_{V}$ and adding positional information as in standard Attention. Next, each of the $Q,K,V\in\mathbb{R}^{L\times d_{\textrm{model}}}$ projections are split into $T=\frac{L}{M}$ chunks of $M$ tokens in each along the sequence dimension, yielding $Q_{i},K_{i},V_{i}\in\mathbb{R}^{M\times d_{\textrm{model}}}$ for $i=1,\ldots,T$. A naive way to reduce the quadratic cost of Attention is to apply it independently on each of these chunks, $A_{i}=\textrm{Attention}(Q_{i},K_{i},V_{i})$ and then concatenate and project them to get the final output, $\textrm{Concatenate}(A_{1},A_{2},\ldots,A_{T})$. However, since the chunks are processed independently, there is no information exchanged between them. Hence, this naive—though efficient—Attention mechanism, which we refer to as “SE-Attn-NoMem”, cannot model time dependencies on contexts larger than the chunk size.
@@ 3.3 Eidetic Retrieval Attention
In this section, we improve upon SE-Attn-NoMem by allowing different chunks to exchange information while minimizing compute; we call this Attention mechanism SE-Attn. To this end, we augment the processing of each chunk with a mechanism to retrieve tokens from previous chunks. In particular, we allow chunk $i$ to retrieve tokens from chunks $11,2,\ldots,i-$ 1 and, when the most relevant chunks are selected, append their tokens to its context (its “expansion span”). SE-Attn takes as input a sequence, $x\in\mathbb{R}^{L\times d}$ and computes projections $Q=xW_{Q}$, $K=xW_{K}$, $V=xW_{V}$ followed by the addition of RoPE [Su et al.(2024)Su, Ahmed, Lu, Pan, Bo, and Liu] embeddings as in standard Attention. As described in Section 3.2, $Q,K,V$ are split into $T$ chunks with $M$ tokens in each, yielding tuples $(Q_{i},K_{i},V_{i})$, $i=1,\ldots,T$. Additionally, $Q,K,V$ are split into a second set of $U$ chunks with $S$ tokens in each, yielding tuples $(Q^{\textrm{Mem}}_{j},K^{\textrm{Mem}}_{j},V^{\textrm{Mem}}_{j})$ where $Q^{\textrm{Mem}}_{j},K^{\textrm{Mem}}_{j},V^{\textrm{Mem}}_{j}\in\mathbb{R}^{S\times d_{\textrm{model}}}$ for $j=1,\ldots,U$. We refer to these tuples as “memory blocks”; in SE-Attn, the query from each chunk, $Q_{i}$, attends not only to $K_{i}$, but also to a set of retrieved memory blocks which populate SE-Attn’s expansion span. In particular, each $Q_{i}$ retrieves $k$ (top-$k$) key/value memory blocks from the past [^2] By “the past,” we mean tokens in the sequence that came before tokens in the chunk.: $\left(K^{\textrm{Mem}}_{\phi_{i}(U)_{1}},V^{\textrm{Mem}}_{\phi_{i}(U)_{1}}\right),\ldots,\left(K^{\textrm{Mem}}_{\phi_{i}(U)_{k}},V^{\textrm{Mem}}_{\phi_{i}(U)_{k}}\right)$ where $\phi_{i}(U)_{j}$ denotes the index of the $j$-th memory block selected by the $i$-th chunk. Retrieved blocks are appended to the chunks’ keys and values, and SE-Attn computes the Attention output for the $i$th chunk as in Equation 3.
@@
$A_{i}^{\textrm{SE-Attn}}$ $=\textrm{Attention}(Q_{i},\tilde{K}_{i},\tilde{V}_{i})$ (3)
Afterwards, each chunk’s Attention outputs are concatenated and projected:
$o^{\textrm{SE-Attn}}$ $=\textrm{Concatenate}(A_{1}^{\textrm{SE-Attn}},A_{2}^{\textrm{SE-Attn}},\dots,A_{T}^{\textrm{SE-Attn}}).$
Note that Equation 3 is a form of Cross-Attention since we are not concatenating memory query tokens to the chunk’s query. This is to preserve causality, as the retrieved tokens cannot attend to the chunk’s tokens.
Memory Retrieval. Each chunk $x_{i}$ must judiciously select which memory blocks to retrieve from the past. To do this efficiently, we associate a 1-dimensional tensor, $c_{j}\in\mathbb{R}^{d_{\textrm{model}}}$, to each of the $j=1,\ldots,U$ memory blocks which act as a compressed representation of each memory block. To determine which memory blocks $Q_{i}$ should attend to, we compute a “relevancy score” between $Q_{i}$ and each $c_{j}$, which measures how relevant memory block $j$ is to chunk $i$. This relevancy score is implemented with a Cross-Attention score between chunks and compressed memory block representations. Recalling that $Q_{i}\in\mathbb{R}^{M\times d_{\textrm{model}}}$, we compute relevancy score $R_{ij}\in\mathbb{R}$ between chunk $i\in\{1,2,\ldots,T\}$ and memory block $j\in\{1,2,\ldots,U\}$ as follows: $R_{ij}=\sum_{t=1}^{M}(Q_{i}c_{j})_{t}.$ $R_{i}\in\mathbb{R}^{U}$ represents the relevancy score between chunk $i$ and all memory blocks. However, since chunk $i$ should only retrieve memory blocks that came before it temporally, we add a mask to $R_{i}$ and then apply Softmax to obtain the final scores, $\tilde{R}_{i}$, between chunk $i$ and all memory blocks as follows:
@@
Compressed Memory Blocks. Next, we discuss how we construct the compressed memory block representations, $c_{j}$. Recently, Landmark Attention [Mohtashami and Jaggi(2023)] considered using “landmark” tokens to obtain compressed representations of memory blocks. We consider using landmark tokens to construct $c_{j}$ in Appendix F. In SE-Attn, we consider a simpler approach which we found to work well. For each memory block, $(Q^{\textrm{Mem}}_{j},K^{\textrm{Mem}}_{j},V^{\textrm{Mem}}_{j})$, we perform standard non-causal Attention, $A^{\textrm{Mem}}_{j}=\mathrm{softmax}\left(\frac{Q^{\textrm{Mem}}_{j}(K^{\textrm{Mem}}_{j})^{T}}{\sqrt{d_{\textrm{model}}}}\right)V^{\textrm{Mem}}_{j}\in\mathbb{R}^{S\times d_{\textrm{model}}}$; we consider non-causal Attention as we are interested in a global representation where all tokens within the memory block can attend to each other. We compute the mean of these weighted memory tokens as the compressed representation: $c_{j}=\frac{1}{S}\sum_{t=1}^{S}(A^{\textrm{Mem}}_{j})_{t}\in\mathbb{R}^{d_{\textrm{model}}}$, where $(A^{\textrm{Mem}}_{j})_{t}$ denotes the $t$-th row of $A^{\textrm{Mem}}_{j}$.
@@ 3.4 Training SE-Attn with LoRA
In this paper, we fine-tune models pre-trained with standard Attention; however, we fine-tune them with SE-Attn, which modifies the pre-trained Attention mechanism by introducing a retrieval mechanism. SE-Attn repurposes the model’s Attention parameters $(W_{Q},W_{K},W_{V},W_{o})$ to perform retrieval. In order to efficiently train the model to learn to use SE-Attn, we use a variant of LoRA [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen]. Recently, [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia] introduced LoRA+ designed to fine-tune Transformer models on long contexts. LoRA fine-tunes Attention parameters with low rank adapter matrices. LoRA+ differs from LoRA by also training embedding and normalization layers.
Recently, [Galim et al.(2024)Galim, Kang, Zeng, Koo, and Lee] found that pure SSMs can be fine-tuned by training the SSM projection layers with LoRA. Since we fine-tune on long contexts, we prioritize efficient training, and consequently apply LoRA only to the Attention layers of our hybrid models. However, it is common for SSM layers to include a 1D convolution layer after their initial projection in order to perform sequence mixing [Zancato et al.(2024)Zancato, Seshadri, Dukler, Golatkar, Shen, Bowman, Trager, Achille, and Soatto, Glorioso et al.(2024b)Glorioso, Anthony, Tokpanov, Whittington, Pilault, Ibrahim, and Millidge, Lieber et al.(2024)Lieber, Lenz, Bata, Cohen, Osin, Dalmedigos, Safahi, Meirom, Belinkov, Shalev-Shwartz, et al., Dao and Gu(2024), De et al.(2024)De, Smith, Fernando, Botev, Cristian-Muraru, Gu, Haroun, Berrada, Chen, Srinivasan, et al.]. The 1D convolution parameters constitute a very small portion of the model parameters ($\sim$0.7for Mamba-2-Hybrid 2.7B), but as discussed further in Section 4.3, we found that training these 1D convolution layers in conjunction with LoRA+ improved our models’ performance on long-context tasks; we refer to this augmented LoRA+ variation as HyLoRA.
@@ 4.1 Experimental Setup
Models. We enhance the recall capabilities of pre-trained hybrid SSM models by fine-tuning them with different Attention layers on spans of tokens longer than the ones used for pre-training. We consider Mamba-2-Hybrid 2.7B [Waleffe et al.(2024)Waleffe, Byeon, Riach, Norick, Korthikanti, Dao, Gu, Hatamizadeh, Singh, Narayanan, et al.] as our representative SSM hybrid model and provide results for Zamba2-Hybrid [Glorioso et al.(2024a)Glorioso, Anthony, Tokpanov, Golubeva, Shyam, Whittington, Pilault, and Millidge] in Appendix D. We also explore expanding the span of Transformer models in Appendix C. These models, obtained from Hugging Face, were pre-trained with a context size of 2048. In our experiments, “Non-fine-tuned” refers to the pre-trained model with no fine-tuning. Similar to [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia], we utilize SE-Attn for efficient fine-tuning and revert to using Full-Attn during evaluation; this is discussed further in Appendix A.
@@
Datasets. We fine-tune models using a common language dataset and provide additional results when fine-tuning on PG-19 [Rae et al.(2019a)Rae, Potapenko, Jayakumar, Hillier, and Lillicrap] and a mixture of language and code dataset in Appendix E.
@@
Baselines. We compare SE-Attn to three Attention variants: standard full Attention (“Full-Attn”) [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin], Sliding Window Attention (“SW-Attn”) [Beltagy et al.(2020)Beltagy, Peters, and Cohan], and Shifted Sparse Attention ($S^{2}$-Attn) [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia]. Full-Attn serves as the paragon, as the other methods aim to approximate it. All Attention implementations use FlashAttention-2 [Dao(2024)].
@@
Training Procedure. We adopt the training recipe used in [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia] to fine-tune our models, with the exception of using HyLoRA instead of LoRA+. See Appendix G for more details.
@@
Evaluation Metrics. We assess the performance of our models across a range of common benchmarks. To assess the predictive capabilities of our fine-tuned models on long sequence lengths, we measure perplexity (PPL) on the PG-19 validation set; however, as discussed further in Appendix B, we do not find PPL to be a faithful measure of a model’s long context abilities. To assess performance on more real-world language tasks in the short and long-context settings, we evaluate our models on various LM Evaluation Harness tasks [Gao et al.(2024)Gao, Tow, Abbasi, Biderman, Black, DiPofi, Foster, Golding, Hsu, Le Noac’h, Li, McDonell, Muennighoff, Ociepa, Phang, Reynolds, Schoelkopf, Skowron, Sutawika, Tang, Thite, Wang, Wang, and Zou]. To measure our models’ performance on long-context tasks, we evaluate on the RULER [Hsieh et al.(2024)Hsieh, Sun, Kriman, Acharya, Rekesh, Jia, Zhang, and Ginsburg] benchmark. We consider eleven RULER tasks, encompassing needle-in-a-haystack, variable-hopping, and aggregation tasks; we aggregate these metrics into five groups as explained in Section I.4. We provide additional long-context benchmarks in Section E.3.
4.2 Results
All of our Mamba-2-Hybrid models are fine-tuned with a context size of 8192. For SE-Attn, we use a block size of 32, and a top-$k$ of 8. For SW-Attn, we use a window size of 4096. $S^{2}$-Attn uses the parameters from [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia]. When fine-tuning Mamba-2-Hybrid using SE-Attn, we found that applying the same chunk sizes at each layer leads to suboptimal downstream performance. Therefore, we segment each sample into chunks of variable sizes (picked randomly from $\{2048,4096\}$). We found this prevents the model from overfitting to a fixed chunk size, which is a hyperparameter chosen a priori, independent of the actual data content. An ablation on SE-Attn with different chunk sizes is provided in Section 4.3.
Perplexity. We provide PG-19 PPL results in Table 1. All fine-tuned models outperform the non-fine-tuned model. SE-Attn yields the closest performance to the paragon model fine-tuned with Full-Attn, outperforming $S^{2}$-Attn and SW-Attn across all context sizes at or above the fine-tuning size.
LM Harness. Next, we evaluate Mamba-2-Hybrid models across short and long-context tasks in the LM Evaluation Harness suite. Our results are provided in Table 1, where we observe that all models perform similarly on short-context tasks—including the non-fine-tuned model—suggesting there is no performance regression when fine-tuning on larger contexts. Furthermore, we observe that fine-tuning with SE-Attn gives the closest performance to fine-tuning with Full-Attn.
Table 1: Fine-tuning Mamba-2-Hybrid with SE-Attn outperforms fine-tuning with $S^{2}$-Attn and SW-Attn on long-context natural language tasks. We evaluate PG-19 validation perplexity (PPL) and observe that fine-tuning with SE-Attn preserves performance at longer contexts better than $S^{2}$-Attn and SW-Attn. On long context tasks from the LM Harness suite, SE-Attn outperforms $S^{2}$-Attn and SW-Attn. See Appendix J for task abbreviation definitions.
RULER. Next, we assess the performance of our models on long-context tasks using the RULER benchmark. LABEL:fig:mamba_ruler_avg_hylora shows the average accuracy across eleven RULER tasks. We observe that fine-tuning with SE-Attn performs similar to fine-tuning with Full-Attn, and outperforms $S^{2}$-Attn and SW-Attn. We provide more detailed RULER results in Figure 6, where we see a substantial improvement on the variable tracking (VT) task, which may be attributed to SE-Attn’s retrieval during fine-tuning.
4.3 Ablations
In this section, we ablate over some of the design choices of SE-Attn on Mamba-2-Hybrid. For this analysis, we consider the RULER benchmark, as it is a strong indicator of performance on long-context tasks. We use the average of the eleven RULER tasks defined in Section I.4.
Does retrieval during training help? In Section 3.2, we introduced SE-Attn-NoMem, a variant of SE-Attn where we do not do any retrieval and process chunks independently. Naturally, we do not expect this to do well due to the lack of shared information across chunks. We confirm this in LABEL:fig:mamba_hydra_ablations(a), where we observe that SE-Attn-NoMem achieves a much lower performance than SE-Attn with retrieval. Furthermore, we also fine-tune with SE-Attn-Random, a variant of SE-Attn that populates its expansion span by retrieving random memory blocks for each chunk. This improves upon SE-Attn-NoMem, indicating that retrieving some information from the past improves performance; however, it does not do as well as SE-Attn, which retrieves blocks based on relevancy.
Chunk size. We found that using a random chunk size during each forward pass improved upon having a fixed chunk size. In LABEL:fig:mamba_hydra_ablations(b), we compare the performance of SE-Attn, which chooses a random chunk size in $\{2048,4096\}$ during each layer’s forward call, to SE-Attn with fixed chunk sizes of 2048 and 4096. Random chunk sizes outperform fixed ones, suggesting a regularizing effect that makes the model more robust to different context lengths.
Block size and top-$k$. For a fixed expansion span size, we might expect that retrieving memory blocks at a finer granularity should perform better than retrieving a smaller number of large blocks; this is because retrieving a larger number of small blocks is as flexible as retrieving a small number of large blocks. However, as illustrated in LABEL:fig:mamba_hydra_ablations(c), we found that fine-tuning SE-Attn with larger block sizes and a smaller top-k gave the best performance. This is possibly due to the increased complexity of the retrieval task, which gets more difficult as the number of possible blocks to retrieve increases. As we are fine-tuning with LoRA, we hypothesize that the model’s capacity may not be sufficient to learn to retrieve effectively. As illustrated in LABEL:fig:mamba_hydra_ablations(d), we found that an expansion span populated with 256 retrieved tokens (32 memory blocks with 8 tokens in each) works best.
HyLoRA for Hybrid Models. We fine-tune our hybrid models using HyLoRA, which builds upon LoRA+ [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia] by also training 1D convolution layers. In LABEL:fig:mamba_ruler_avg_hylora(b) we show that, when evaluated on RULER tasks, fine-tuning SE-Attn models with HyLoRA gives the strongest performance. In Appendix B, we study the effect of the LoRA rank and apply HyLoRA to Full-Attn.
4.4 Empirical Runtime Analysis
While hardware-efficient Attention implementations, such as FlashAttention [Dao(2024)], enable linear memory scaling, computational complexity remains quadratic in sequence length. In LABEL:fig:mamba_fwd_bwd_anal(a), we profile different Attention layers on various context sizes for one training step and see that SE-Attn is much faster than $S^{2}$-Attn and Full-Attn. Moreover, the runtime of SE-Attn is similar to that of SW-Attn, despite SE-Attn’s much longer Attention span. We also note that the runtime increase of Full-Attn/FlashAttention compared to SE-Attn is 4-5$\times$ for long sequences, while the memory peak increase is 17%, as shown in LABEL:fig:mamba_fwd_bwd_anal(b). SE-Attn has a minor memory overhead because it requires storing partial block summary statistics in order to retrieve tokens from the past. However, this overhead is outweighed by its faster runtime. As depicted in LABEL:fig:mamba_fwd_bwd_anal(c), our method has a strong runtime-memory trade-off—similar to that of SW-Attn, but with the added benefit of improved performance on language benchmarks. See Appendix H for a theoretical runtime analysis.
5 Conclusion
We close with the limitations of our method. Although we have conducted experiments at a relatively large model scale (2.7B) and long contexts (up to 32k), testing our method for efficiently adapting larger models on longer contexts is paramount. Furthermore, while we utilize datasets that require modeling of long-range dependencies, we found that perplexity-based tasks do not faithfully measure models’ capabilities to handle long contexts, and instead tasks like RULER provide better signals of long-context capabilities. However, RULER is mostly a synthetic dataset and does not cover more nuanced tasks that require reasoning over long documents. Validating our method on more complex long-range benchmarks is a promising area for future work.
Appendix A Evaluating with Full Attention